OpenStudy (anonymous):

The square root of 3 is 1.73. Find the square root of 147? (answer is 12.11 but I don't know how to work ...

OpenStudy (cj49):

147 = 49 *3 = sqrt(49) sqrt(3) final ans 7 sqrt3 7*1.73

OpenStudy (anonymous):

\[√147= √(3 * 49) \] \[(√3) (√49)= (1.73) (7)= 12.11\]

OpenStudy (anonymous):

thank you all